OpenStudy (anonymous):

If gas particles start colliding with the walls of their metallic container with increased force, what is their direct effect? lower gas pressure higher gas pressure lower volume of gas higher volume of gas

OpenStudy (aaronq):

A metal container will not expand, so the volume will not change. This eliminates 2 of the options. \(\sf Pressure=\dfrac{Force}{Area}\) Say the force initially was 100 N, and it increased to 200 N. \(\sf Pressure=\dfrac{100~N}{1~m^2}=100~Pa\) \(\sf Pressure=\dfrac{200~N}{1~m^2}=200~Pa\) An increased of force in the collisions with the container walls increased the pressure.